OpenStudy (anonymous):

Can someone please help me? A rectangular lot has a perimeter of 350 feet. An expensive fencing along the lot's lengths costs $27.60 per foot. An inexpensive fencing along the widths costs $4.70 per foot. The total cost of the fencing comes to $4,622. What are the lot's dimensions?

OpenStudy (anonymous):

You need to set up the two equations. Let L be length and W be width. Perimeter = 350 ft. So, what is that in terms of L and W?

OpenStudy (anonymous):

2L+2w=350 ?

OpenStudy (anonymous):

YES!!

OpenStudy (anonymous):

Now, the second equation. Cheap fence using along widths and expensive used along widths. What is the cost equation in terms of L and W?

OpenStudy (anonymous):

2(27.60)+2(4.70)=4622

OpenStudy (anonymous):

Well....its is 2L(27.60) + 2W(4.70) = 4622. => 55.20L + 9.40W = 4622 So, now you have two equations in two variables. Solve them using substitution or elimination.

OpenStudy (anonymous):

sweet thanks!